# Robust Deep Network Learning of Nonlinear Regression Tasks by Parametric Leaky Exponential Linear Units (LELUs) and a Diffusion Metric

### v1

@@ Abstract
This document proposes a parametric activation function (ac.f.) aimed at improving multidimensional nonlinear data regression. It is a established knowledge that nonlinear ac.f.’s are required for learning nonlinear datasets. This work shows that smoothness and gradient properties of the ac.f. further impact the performance of large neural networks in terms of overfitting and sensitivity to model parameters. Smooth but vanishing-gradient ac.f.’s such as ELU or SiLU have limited performance and non-smooth ac.f.’s such as RELU and Leaky-RELU further impart discontinuity in the trained model. Improved performance is demonstrated with a smooth “Leaky Exponential Linear Unit”, with non-zero gradient that can be trained. A novel diffusion-loss metric is also proposed to gauge the performance of the trained models in terms of overfitting.
@@ 1 Introduction
It is typical in engineering applications that a modelled system is multidimensional, i.e., a representative output parameter of the system depends on many system operational variables. The model might also be structured following a mesh-like combination of the system variables, which were input on a computational model or test bench for the data generation. In the example to be addressed in this document, a metric of an electric motor efficiency is dependent on three variables, representative of the motor speed, power, and electric feed. This seemingly simple problem has proven to be surprisingly difficult to be trained for data regression. In an effort on Gaussian Process Regression (GPR) methods [1], the author developed a loss based on a diffusion operator to detect and avoid overfitting resulting from the highly nonlinear training landscape.
@@
The scientific and engineering communities are interested in modelling techniques that are robust to parameters and that provide good generalisation. The selection of the activation function (ac.f.) of a NN heavily impacts such objectives, and it is known that nonlinear ac.f.’s are necessary to be able to model nonlinear datasets [3]. However, most of the literature focuses on large models for classification problems [4, 5, 6, 7], and regression problems are typically small [8] and the concern is usually with underfitting.
@@
A dense feedforward NN was set up to learn the aforementioned motor model, for which relatively large depths and widths were required. In general, the “global” overfitting patterns of the GPR methods seen in [1] are not observed with the NN. As shown in Sect. 4.2, rather “localised” low-amplitude small-lengthscale features are observed – in this sense, “localised” refers to over-expression between adjacent data points, thus not a “global” long wavelength pattern spanning through many data points. In special, the regions where two of the variables are fast changing but the third one is not, seemed to exacerbate the issue.
@@
A motivation to pursue a deeper understanding of the observed issues comes from this remark in [9], that “a priori knowledge of a particular task into a learning algorithm (…) generally improves performance, if the incorporated knowledge is relevant to the task.” The author proposes an ad hoc metric of the flexibility of the ac.f. as a ratio of its maximum and minimum gradients. In considering this metric, it can be observed that highly flexible ac.f.’s are prone to over-expressing in large models, whereas the conventional linear unit has zero flexibility and thus only represents linear functions (a common knowledge [3]). The current work proposes a parametric, trainable flexibility that enhances the a priori “fitness” of an ac.f. for more generalisable nonlinear learnings.
@@
Furthermore, concern with ac.f. smoothness has motivated other work [5, 7], and the notion of trainable/adaptive coding in ac.f. has demonstrated successful behaviour [10, 11, 12]. This outcome is herein achieved with a simple ac.f. definition, termed Leaky Exponential Linear Unit (Leaky ELU, or LELU). Importantly, the focus here is broader in the sense that it focuses on a set of behavioural features of an ac.f. for generalisation, one of them being the discussed flexibility, but also robustness (i.e., reduced dependency) with model size. Quoting [12] “there is no rule of thumb of choosing an optimal activation function.” Essentially, we want to decrease the effort in finding an optimal ac.f. by providing an adaptive framework that is able to code the degree of presence of particular phenomena in the input in a smooth and ever proportional way.
@@
The localised overfitting behaviour is demonstrated on canonical unidimensional problems, with model sizes representative of that for the motor model. It is shown that classical ac.f.’s have mixed performance in the problems whereas the proposed LELU is robust to model size. The observed properties of the LELU are further demonstrated in the motor model.
@@ 2 Activation Functions
This scores is zero for inflexible ac.f.’s, such as the Linear Unit (LU), which can only represent linear functions. On the other extreme, an ac.f. with score 1 is highly flexible. Table 1 shows the score for the ac.f.’s plotted in Fig. 1.
Table 1: Flexibility score of classical activation functions by application of Eq. 1. The parametrized activation functions affect the flexibility score by their control parameter.
Activation $\eta(\phi)$
@@
In the current work, it is shown that for nonlinear regressions tasks the following features of an ac.f.’s negatively impact the model performance:
@@
where the trainable parameter $\beta$ controls the slope of the activation function in the negative region (see Fig. 1).
@@
These requirements are imposed onto the ac.f. construction to enhance the performance for nonlinear regressions tasks. The regression results show that this set of features inherently produces a behaviour similar to batch normalisation without the need for regularisation, as for instance explicitly required in [7, 16].
@@ 3.1 Neural Network Configuration
The NN is implemented in Python using the Keras library [17] and trained to minimise the mean absolute error (MAE) employing the Adam optimizer [18]. The minimisation of the MAE has provided better results for the regression tasks when compared to a mean squared error (MSE). The learning rate is step-annealed based on the MAE loss – more details are later provided for each model. Neurons are initialised with a HeNormal distribution [4].
Regularization in regression tasks is essentially a trade-off: it helps prevent overfitting and can improve training stability, but by definition the minimum achievable training error is limited by the strength of the regularization. Therefore, no regularisation is used as default in the current work, nonetheless this aspect is also herein briefly assessed since it is a common practice.
The use of a custom diffusion operator as a testing metric of performance allows for training on the full dataset. This metric is assessed via its MSE and described in the following subsection.
@@ 3.2 Diffusion-Loss Metric
The diffusion-loss formulation herein applied follows the more complete description in [1]. This methodology allows for the creation of a hierarchical testing metric from the training data field without requiring splitting a training dataset into testing sets. Training is performed on the complete available dataset and offers best chance for generalisation.
@@
~$\tilde{\nabla}y_{i}=\frac{1}{3~{}(\Delta/2)^{2}}\frac{\big{|}\hat{y}_{i+1}-\hat{y}_{i+\frac{1}{2}}-\hat{y}_{i-\frac{1}{2}}+\hat{y}_{i-1}\big{|}}{\hat{y}_{i+1}+\hat{y}_{i+\frac{1}{2}}+\hat{y}_{i-\frac{1}{2}}+\hat{y}_{i-1}},$ (4)
@@
~$\tilde{\nabla}y_{i,j,k}^{p}=\frac{1}{3~{}(\Delta/2)^{2}}\frac{\big{|}y_{i+1,j+1,k+1}-y_{i+\frac{1}{2},j+\frac{1}{2},k+\frac{1}{2}}-y_{i-\frac{1}{2},j-\frac{1}{2},k-\frac{1}{2}}+y_{i-1,j-1,k-1}\big{|}}{y_{i+1,j+1,k+1}+y_{i+\frac{1}{2},j+\frac{1}{2},k+\frac{1}{2}}+y_{i-\frac{1}{2},j-\frac{1}{2},k-\frac{1}{2}}+y_{i-1,j-1,k-1}},$ (5)
@@
Note that the diffusion sensors are only computed at the interior of the original mesh, where the complete stencil is available. The reader is referred to [1] for further and more general discussions and information about the methodology as well as snippets of code implementation.
@@ 4 Results and Discussions
We are interested in multidimensional problems that typically require large models to capture all their features. The author argues that using canonical problems with isolated features and fitting small NNs to them do not guarantee generalisation of the approach for those large, real-world problems. We want to be able code the degree of presence of series of particular phenomena that might compose large problems. Therefore, in this work, we take large NNs that are required to represent large, complex models, and apply them to canonical problems in search of ac.f.’s that offer robustness whilst coding the degree of presence of different and composed phenomena. The goal is to provide an ac.f. that works well regardless of model size and that can scale effectively from small to large networks, thereby enabling potentially more robust and less time consuming regression tasks.
@@ 4.1 Unidimensional Problems
The proposed unidimensional problems [^1] Source files available in http://github.com/ebiga/nn_1d_overfit_ex represent certain important features of the target motor model. A hyperbolic tangent function is chosen for its non-monotonic change of slope, whereas an exponential function is chosen for its fast change of properties. Numerical experiments with a NN of depths $n=[7,8]$ and widths $m=[120,240]$ (fixed per layer) are executed for the ac.f.’s listed in Table 2.
@@ 4.1.1 Hyperbolic Tangent Function
A 1d equation given by $f(x)=0.5+0.5\tanh(5x),~{}x\in[-1,1]$ is used to generate two sets of equally spaced, noiseless $(7,14)$ training points. The model is trained for 15000 epochs with batch size 3. The learning rate is scheduled in the range $[1e-3,1e-6]$ with factor 0.5, patience 500, and cooldown 100.
Table 2: Losses of the trained models for the hyperbolic tangent data.
Activation Training MAE ×$\times 10^{-6}$ Diffusion MSE ×$\times 10^{-3}$
$14pts$ $7pts$ $14pts$ $7pts$
Leaky ReLU $(\alpha=0.2)$ 14 10 90 216
LELU $(\beta=0.3)$ 20 27 16 25
LELU $(\beta=0.4)$ 22 101 49 72
LELU $(\beta=0.6)$ 54 68 15 73
Softplus 71 21 16 220
SiLU 23 19 13 49
tanh 17 8 156 372
ELU 17 88 55 234
ReLU 8 5 125 241
Table 2 presents the training losses and diffusion metrics of the trained model. The MAEs are generally low and care must be taken that the models are not tending to memorise the training points, the performance of the models is therefore truly assessed by the diffusion metric. The SiLU and LELU ($\beta=0.3$) achieve low MAE and low diffusion losses, and the performance is not strongly lost with the reduction of the number of training points. This result is indicative of less overfitting, which is corroborated by the function plots in Fig. 5. Strongly overfitting is observed with some ac.f.’s in this setup, whereas the proposed LELU is relatively robust to the change in the number of training points. Not shown for the sake of conciseness, very similar results are observed with a $7\times 120$ model as well.
@@
Figure 5: Trained hyperbolic tangent function plots for the different activation functions.
@@ 4.1.2 Exponential Function
A 1d exponential equation given by $f(x)=0.1^{x},~{}x\in[-1,1]$ is used to generate a set of $12$ equally spaced, noiseless training data points. The model is trained for 20000 epochs with batch size 3. The learning rate is scheduled in the range $[1e-3,2e-7]$ with factor 0.5, patience 500, and cooldown 100.
Table 3: Losses of the trained models for the exponential function data.
Activation Training MAE ×$\times 10^{-6}$ Diffusion MSE ×$\times 10^{-3}$
$n=8$ $n=7$ $n=8$ $n=7$
Leaky ReLU $(\alpha=0.2)$ 4 31 12 16
LELU $(\beta=0.3)$ 50 19 3 2
LELU $(\beta=0.4)$ 24 33 2 2
LELU $(\beta=0.6)$ 34 62 1 1
Softplus 120 133 5 2
SiLU 7 25 13 14
tanh 21 14 28 37
ELU 18 12 18 20
ReLU 4 5 25 19
Table 3 presents the training losses and diffusion metrics of the trained model. As in the previous case, the lower the MAE, the higher is the tendency of the model to memorise the training data, with a resulting increase in the diffusion metric. For this case, both SiLU and Softplus show a poorer performance and trade the training and diffusion (overfitting) errors. The LELUs are able to adequately meet both low MAE and diffusion errors. The trained model predictions are plotted in Fig. 6. As in the previous case, different levels of overfitting are observed with the classical ac.f.’s in this setup, whereas generally robust results are obtained with the LELU.

### v2

@@ Abstract
This document proposes a parametric activation function (ac.f.) aimed at improving multidimensional nonlinear data regression. It is a established knowledge that nonlinear ac.fs. are required for learning nonlinear datasets. This work shows that smoothness and gradient properties of the ac.f. further impact the performance of large neural networks in terms of overfitting and sensitivity to model parameters. Smooth but vanishing-gradient ac.fs. such as ELU or SiLU (Swish) have limited performance and non-smooth ac.fs. such as RELU and Leaky-RELU further impart discontinuity in the trained model. Improved performance is demonstrated with a smooth “Leaky Exponential Linear Unit”, with non-zero gradient that can be trained. A novel diffusion-loss metric is also proposed to gauge the performance of the trained models in terms of overfitting.
@@ 1 Introduction
1.1 Perspectives and Motivation
It is typical in engineering applications that a modelled system is multidimensional, i.e., a representative output parameter of the system depends on many system operational variables. The model might also be structured following a mesh-like combination of the system variables, which were input on a computational model or a test bench for the data generation. In the example to be addressed in this document, a metric of an electric motor efficiency is dependent on three variables, representative of the motor speed, power, and electric feed. This seemingly simple problem has proven to be surprisingly difficult to be trained for data regression. In an effort with Gaussian Process Regression (GPR) [1], the author developed a loss based on a diffusion operator to detect and avoid overfitting resulting from the highly nonlinear training landscape.
@@ 1.1 Perspectives and Motivation
The scientific and engineering communities are interested in modelling techniques that are robust to parameters and that provide good generalisation. The selection of the activation function (ac.f.) of a NN heavily impacts such objectives, and it is known that nonlinear ac.fs. are necessary to model nonlinear datasets [3]. Furthermore, smoothness (e.g., the classical ELU [4]) as well as trainable/adaptive coding properties for ac.fs. receive consistent attention in the literature, and the reader is referred to the thorough summary in [5] for further information. More complex approaches such as ac.fs. based on GPRs [6], rational functions [7], Kronecker Neural Networks [8], or even on model structural changes [9], have also been proposed. Of importance to the current work, several trainable variants are available solely for ReLU and ELU (again summarised in [5]).
The choice of an ac.f. can in fact be overwhelming and other broad-scope reviews are available in [10, 11]. However, the literature generally focuses on large models for classification problems whereas regression problems are typically small. Even though such reviews provide interesting background for further exploration, regression tasks can be more demanding than classification since they require continuity. An interesting example of such difficulties is a compromise termed “benign overfitting” [12], in which localised spikes near noisy training points are accepted for a model that generalises well elsewhere.
@@
In the current work, a dense feedforward NN is developed to learn the aforementioned motor model. Relatively large depths and widths are required to properly learn such nonlinear problem [13, 14]. As shown in Sect. 4.2, over-expression between adjacent data points is observed with classical ac.fs. and the use of regularisation to avoid such patterns incur in larger training errors. In special, the regions where two of the variables are fast changing but the third one is not, seemed to exacerbate the issue.
@@
A motivation to pursue a deeper understanding of the observed issues comes from this remark in [15], that “a priori knowledge of a particular task into a learning algorithm (…) generally improves performance, if the incorporated knowledge is relevant to the task.” The author proposes an ad hoc metric of the flexibility of the ac.f. as a ratio of its maximum and minimum gradients. In considering this metric, it can be observed that highly flexible ac.fs. are prone to over-expressing in large models, whereas the conventional linear unit has zero flexibility and thus only represents linear functions (a common knowledge [3]).
1.2 Contributions of the Work
@@
The current work proposes a parametric, trainable flexibility that enhances the a priori “fitness” of an ac.f. for more generalisable nonlinear learnings. This outcome is herein achieved with a simple ac.f. definition, termed Leaky Exponential Linear Unit (Leaky ELU, or LELU). Importantly, the focus here is broader in the sense that it focuses on a set of behavioural features of an ac.f. for generalisation, one of them being the discussed flexibility, but also robustness (i.e., reduced dependency) with model size. Quoting [5] “there is no rule of thumb of choosing an optimal activation function.” Essentially, we want to decrease the effort in finding an optimal ac.f. by providing an adaptive framework that is able to code the degree of presence of particular phenomena in the input in a smooth and ever proportional way.
@@
1.3 Organisation
The rationale for the ac.f. definitions are initially discussed. The NN setup and diffusion loss definitions are then presented. The localised overfitting behaviour is demonstrated on canonical unidimensional problems, with model sizes representative of that for the motor model. It is shown that classical ac.fs. have mixed performance in the problems whereas the proposed LELU is robust to model size. The observed properties of the LELU are further demonstrated in the motor model.
@@ 2 Activation Functions
This scores is zero for inflexible ac.fs., such as the Linear Unit (LU), which can only represent linear functions. On the other extreme, an ac.f. with score 1 is highly flexible. Table 1 shows the score for the ac.fs. plotted in Fig. 1, and includes a wide range of original to more modern ac.fs.
Table 1: Flexibility score of several activation functions by application of Eq. 1. The parametrizable activation functions can affect their flexibility score (examples in Fig. 1).
Activation $\mathbf{\phi}$ $\eta(\phi)$
@@
  • 1.
    *The term SiLU instead of Swish [20] is adopted for consistency with the Keras library. For the current purposes, the other modern option Mish [21] showed very similar performance to SiLU/Swish.
@@
In the current work, it is shown that for nonlinear regressions tasks the following features of an ac.f. negatively impact the model performance:
@@
where the trainable parameter $\beta$ controls the slope of the activation function in the negative region (see Fig. 1). Interestingly, the elastic ELU in [22] proposes a change in the slope of the ac.f. however on the positive, right-hand side.
@@
These requirements are imposed onto the ac.f. construction to enhance the performance for nonlinear regressions tasks. The regression results show that this set of features inherently produces a behaviour similar to batch normalisation without the need for regularisation, as for instance explicitly remarked in [23, 7, 24].
@@ 3.1 Neural Network Configuration
The NN is implemented in Python using the Keras library [25] and trained to minimise the mean absolute error (MAE) employing the Adam optimizer [26], as summarised in Algorithm 1.
Algorithm 1 Training procedure of the artificial neural network.
1:Training data $(\vec{x},f(\vec{x}))$. $\triangleright$ From system models or generated from canonical functions.
2:Trained neural network and diffusion-loss metric.
3:procedure NN Setup
4:  Define the NN architecture (depth and width, see Fig. 2).
5:  Configure learning rate annealing and number of epochs for the Adam optimizer.
6:  Set activation functions. $\triangleright$ Regularisation applied when intended.
7:  Initialize weights with chosen initializer.
8:end procedure
9:while max epoch not reached do
[…]
11:end while
12:procedure Diffusion metric $\triangleright$ Detailed in Sect. 3.2.
13:  Make predictions at staggered mesh nodes and at training points.
14:  Compute diffusion terms at the training points according to Eq. 7.
15:  Compute the diffusion loss applying Eq. 8.
16:end procedure
The minimisation of the MAE has provided better results for the regression tasks when compared to a mean squared error (MSE). The learning rate is step-annealed based on the MAE loss and the weights are initialised by proper techniques, with more details later provided for each training case.
Regularization in regression tasks is essentially a trade-off: it helps prevent overfitting and can improve training stability, but by definition the minimum achievable training error is limited by the strength of the regularization. Therefore, no regularisation is used as default in the current work, nonetheless this aspect is also herein assessed since it is a common practice.
@@ 3.2 Diffusion-Loss Metric
The use of the diffusion loss in Algorithm 1 provides an independent testing metric of performance and also allows for training on the full dataset, as discussed in the current subsection. The diffusion-loss formulation herein applied follows the more complete description in [1]. This methodology allows for the creation of a hierarchical testing metric from the training data field without requiring splitting a training dataset into testing sets. Training is performed on the complete available dataset and offers best chance for generalisation.
@@
$\tilde{\nabla}y_{i}=\frac{1}{3~(\Delta/2)^{2}}\frac{\big|\hat{y}_{i+1}-\hat{y}_{i+\frac{1}{2}}-\hat{y}_{i-\frac{1}{2}}+\hat{y}_{i-1}\big|}{\hat{y}_{i+1}+\hat{y}_{i+\frac{1}{2}}+\hat{y}_{i-\frac{1}{2}}+\hat{y}_{i-1}},$ (4)
@@
$\tilde{\nabla}y_{i,j,k}^{p}=\frac{1}{3~(\Delta/2)^{2}}\frac{\big|y_{i+1,j+1,k+1}-y_{i+\frac{1}{2},j+\frac{1}{2},k+\frac{1}{2}}-y_{i-\frac{1}{2},j-\frac{1}{2},k-\frac{1}{2}}+y_{i-1,j-1,k-1}\big|}{y_{i+1,j+1,k+1}+y_{i+\frac{1}{2},j+\frac{1}{2},k+\frac{1}{2}}+y_{i-\frac{1}{2},j-\frac{1}{2},k-\frac{1}{2}}+y_{i-1,j-1,k-1}},$ (5)
@@
Note that the diffusion sensors are only computed at the interior of the original mesh, where the complete stencil is available. The loss of the diffusion sensor is finally computed as the MSE by
$\mathrm{MSE}_{diffusion}=\frac{1}{E}\sum_{e=1}^{E}(\nabla y_{e}-\tilde{\nabla}y_{e})^{2},$ (8)
where $E$ is the internal nodes of the training mesh. The reader is referred to [1] for further and more general discussions and information about the methodology as well as snippets of code implementation.
@@ 4 Results and Discussions
We are interested in multidimensional problems that typically require large models to capture all their features. The author argues that using canonical problems with isolated features and fitting small NNs to them do not guarantee generalisation of the approach for those large, real-world problems. We want to be able to code the degree of presence of series of particular phenomena that might compose large problems. Therefore, in this work, we take large NNs that are required to represent large, complex models, and apply them to canonical problems in search of ac.fs. that offer robustness whilst coding the degree of presence of different and composed phenomena. In essence, the context within is models with excess capacity surpassing network size guidelines [27] to better fit regions of high non-linearity [14] and that likely exceed the sample size [13, 28]. The goal is to provide an ac.f. that works well regardless of model size and that can scale effectively from small to large networks, thereby enabling potentially more robust and less time consuming regression tasks.
@@ 4.1 Unidimensional Problems
The proposed unidimensional problems [^1] Source files available in http://github.com/ebiga/nn_1d_overfit_ex represent certain important features of the target motor model. A hyperbolic tangent function is chosen for its non-monotonic change of slope, whereas an exponential function is chosen for its large and fast change of properties. Numerical experiments with NNs of different depths and widths are executed with the previously presented ac.fs. A set 4 seeds of He-Normal [18] and 2 seeds of Glorot-Uniform [29] weight initialisers have been generated and the results are presented in the form of mean and standard deviation. Where applicable, experiments with L2 regularisation are also reported.
@@ 4.1.1 Hyperbolic Tangent Function
A 1d equation given by $f(x)=0.5+0.5\tanh(5x),~x\in[-1,1]$ is used to generate two noiseless sets of equally spaced $14$- and a more challenging $7$-point cases. The models are trained with $(n\times m)=(7\times 120)$ and $(8\times 240)$, for 15000 epochs with batch size 3. The learning rate is scheduled in the range [1e3,1e6][1e-3,1e-6] with factor 0.5, patience 500, and cooldown 100.
Table 2: Mean and standard deviation of the losses of different model setups for the hyperbolic tangent with 1414 training points.
Activation Training MAE $\times 10^{-6}$ Diffusion MSE $\times 10^{-2}$
$n=7$ $n=8$ $n=7$ $n=8$
ELU (15;7) (18;8) (39;14) (80;98)
Leaky ReLU $(\alpha=0.2)$ (10;3) (15;5) (126;56) (156;59)
LELU $(\beta=0.3)$ (19;12) (27;33) (16;4) (25;4)
LELU $(\beta=0.4)$ (19;10) (30;99) (18;3) (22;6)
LELU $(\beta=0.6)$ (22;34) (63;57) (14;3) (18;5)
ReLU (7;3) (7;3) (216;59) (275;34)
ReLU-L2 (5;15) (19;23) (107;3) (108;7)
SiLU (10;3) (12;8) (17;0) (17;0)
SiLU-L2 (171;118) (136;89) (17;2) (17;2)
Softplus (43;14) (65;15) (18;0) (36;14)
Tanh (37;39) (69;28) (249;126) (272;528)
Table 3: Mean and standard deviation of the losses of different model setups for the hyperbolic tangent with 77 training points.
Activation Training MAE $\times 10^{-6}$ Diffusion MSE $\times 10^{-2}$
$n=7$ $n=8$ $n=7$ $n=8$
ELU (27;12) (22;32) (41;8) (54;50)
Leaky ReLU $(\alpha=0.2)$ (7;7) (10;6) (37;38) (54;39)
LELU $(\beta=0.3)$ (17;14) (57;54) (14;4) (28;10)
LELU $(\beta=0.4)$ (40;15) (60;36) (8;4) (9;2)
LELU $(\beta=0.6)$ (35;24) (64;33) (14;6) (16;5)
ReLU (6;5) (6;4) (37;63) (53;115)
ReLU-L2 (8;5) (2;2) (72;4) (74;4)
SiLU (9;5) (10;5) (5;4) (7;3)
SiLU-L2 (5;106) (170;142) (9;1) (8;0)
Softplus (21;11) (51;28) (58;35) (71;35)
Tanh (13;8) (10;22) (96;42) (86;39)
Tables 2 and 3 present the training and diffusion [^2] In this case for which the function values approach zero, the denominators in Eqs. 3 and 4 are not implemented. losses for $14$ and $7$ training points, respectively. As expected, the models tend to perform better with the larger number of training points, however the errors are not necessarily much smaller. The models on the lower end of MAEs are generally tending to memorise the training points, as seen in the function plots in Fig. 5, so the performance of the models is fundamentally gauged by the diffusion metric.
SiLU and Softplus present adequate diffusion metrics for this hyperbolic tangent function, closely followed by LELU. Softplus however loses performance with the deeper model and with the lower number of points. All other ac.fs. perform rather poorly, and the application of regularisation (strength 1e41e-4) does not improve or even worsens the performance. For this particular problem, LELU can perform closely to SiLU at the expense of higher training MAE. It is expected that the optimal parameter $\beta$ of LELU is problem and model dependent, which can be observed in the presented results, and that is an intended property.
@@
Figure 5: Trained hyperbolic tangent function plots for the different activation functions for the He-Normal initialisation (no seed).
@@ 4.1.2 Exponential Function
A 1d exponential equation given by $f(x)=0.1^{x},~x\in[-1,1]$ is used to generate a set of $12$ equally spaced, noiseless training data points. The models are trained with $(n\times m)=(7\times 240)$ and $(8\times 240)$, for 15000 epochs with batch size 3. The learning rate is scheduled in the range [1e3,2e7][1e-3,2e-7] with factor 0.5, patience 500, and cooldown 100.
Table 4: Mean and standard deviation of the losses of different model setups for the exponential function with 1212 training points.
Activation Training MAE $\times 10^{-6}$ Diffusion MSE $\times 10^{-3}$
$n=7$ $n=8$ $n=7$ $n=8$
ELU (21;62) (20;15) (12;9) (19;16)
Leaky ReLU $(\alpha=0.2)$ (10;9) (11;14) (14;4) (15;3)
LELU $(\beta=0.3)$ (20;6) (36;25) (2;1) (3;2)
LELU $(\beta=0.4)$ (44;57) (40;99) (2;1) (2;1)
LELU $(\beta=0.6)$ (38;227) (52;184) (2;0) (1;0)
ReLU (8;6) (7;8) (21;4) (18;4)
ReLU-L2 (6;1) (7;2) (18;8) (24;7)
SiLU (17;95) (12;75) (15;8) (13;7)
SiLU-L2 (155;683) (104;567) (8;46) (7;5)
Softplus (115;111) (97;166) (2;1) (4;0)
Tanh (11;5) (157;165) (33;9) (24;23)
Table 4 presents the training and diffusion losses of the trained models. As in the previous case, the lower the MAE, the higher is the tendency of the model to memorise the training data with a resulting increase in the diffusion metric (see Fig. 6). Different levels of overfitting are observed with the classical ac.fs. (including regularisation) in this setup, whereas generally robust results are obtained with LELU. Contrary to the previous observations with the hyperbolic tangent results, both SiLU and Softplus show a poorer performance in terms of training MAE or diffusion MSE. The LELUs are able to adequately meet both low MAE and diffusion errors, with the tendency for slightly better overall performance with lower $\beta$ due to lower MAE.